# AAPM: Large Language Model Agent-based Asset Pricing Models

### v1

@@ Appendix A Hyperparam Search
We employed PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) for parallel training.
Appendix B Dataset visualizations
Figure 7: The word cloud of topics of the WSJ business news over time in our dataset (Bottom) compared to the corresponding risk adjusted market return (Top).
Figure 6 illustrates the variations in the number of articles and assets over time. We analyze the primary topics discussed in the news articles within our dataset across different periods. The topics were determined based on the titles of the news articles for each season. Common words such as "US," "Stock," and "Market" were excluded as they did not effectively represent the event’s topic. The resulting word cloud is shown in Figure 7. It is clear that the economy is mainly influenced by various long-term events. It begins with a gradual decline in the emphasis on COVID. Then, the focus shifted towards managing inflation and the decisions made by the FED. The banking crisis at the start of 2023 soon became the new central point, followed by the acknowledgment of AI as a key driver for the economy, mainly due to the success of LLMs. This indicates that these event trends have the potential to serve as strong predictors of economic indicators and the market. This is also reflected in Appendix C, where we evaluated that news articles have significant predictive power for economic indicators and market trends.
Figure 8: The most frequent mentioned stock tickers in the news.
{beautifulbox}
Appendix C News as Financial Economic Predictor
To explore the predictive capability of business news on financial and economic dynamics, we conduct an experiment using refined news features to forecast the economic indicators in Appendix C.1 and market movements in Appendix C.2. We embed the refined news directly and use the daily averaged embeddings of the refined news as predictors in our experiments.
C.1 Economic Indicators
We assess the capability of news features to forecast the daily percentage changes in typical and most popular macroeconomic indicators in different topics sourced from the FRED database [^3] https://fred.stlouisfed.org/. These indicators encompass the stock market (SP500), the market yield on U.S. Treasury Securities at a 10-Year constant maturity (DGS10), Moody’s seasoned Baa corporate bond minus the federal funds rate (BAAFF), the 10-year breakeven inflation rate (T10YIE), Brent crude oil prices (DCOILBRENTEU), and the 30-year fixed-rate conforming mortgage index (OBMMIC30YF). The findings are illustrated in Figure 9. The forecasted results exhibit a high degree of accuracy, as evidenced by the high R2 score. This implies that news provides valuable insights for predicting macroeconomic indicators.
C.2 Stock Price Predictor
We further investigate the predictive power of news features to the price movements of individual stocks. We chose 8 typical stocks that has been frequently mentioned in the new from our analysis in Appendix 8, and used refined news features as predictors to estimate their daily percentage price changes. The results are displayed in Figure 10, with the corresponding R2 scores given in brackets. We note high accuracy and R2 scores for all selected stocks, suggesting that news can significantly help in forecasting stock prices. However, it is important to recognize that due to non-stationarity and the risk of overfitting, stock price prediction cannot be directly applicable as asset pricing Kelly et al. (2023), but it provides insights into the value of the news in the pricing of individual stock.
Appendix D Illustration of the Ablation Baselines
@@
Figure 11: Example of the Hybrid agent baseline analyze a raw news without iterative refinement of analysis report as well as the macroeconomic and market trend notes.
@@ E.2 Iterative analysis
You are a helpful assistant designed to analyze the business news to assist portfolio management. You will help me analyze this latest news from The Wall Street Journal and provide an analysis report, then I will search the relevant news or articles from the knowledge base based on your analysis report to help you refine it iteratively in multiple rounds. Let’s start with this latest news, provide your analysis report, and I will help you refine it with the relevant information later, if you think this news is completely not helpful for investment now or future, call skip function to skip it, do not skip it if it may contain helpful information to future investment:
@@
Based on your current analysis report, I found those potentially relevant news and excerpts from the knowledge base, please refine your analysis report with this information:
@@
Based on your current analysis report, I found those potentially relevant news and excerpts from the knowledge base, now finish your analysis report with them:
@@ E.3 Update of note
Now, given the latest news and the analysis report, update the macroeconomic summary with the new insights and impacts from the news. Include any relevant information that could influence the global economic outlook, such as geopolitical events, policy changes, or economic indicators. You should also take note of any important notes about investment trend and chances. Here are the latest news and the analysis report:

### v2

@@ Appendix A Hyperparam Search
We employed PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) for parallel training.
Appendix B Illustration of the Ablation Baselines
@@
Figure 5: Example of the Hybrid agent baseline that analyzes raw news without iterative refinement of analysis report as well as the macroeconomic and market trend notes.
Appendix C Analysis on Dataset
In this section, we analyze the dataset with a focus on the first two years, a highly complicated period as illustrated in Section 3 to provide better insights on the news serves as market signals.
Figure 7: The word cloud of topics of the WSJ business news over the first two years in our dataset (Bottom) compared to the corresponding risk-adjusted market return (Top).
News and Topics over time.
Figure 6 illustrates the variations in the number of articles and assets over time for the first two years. We analyze the primary topics discussed in the news articles within our dataset across different periods. The topics were determined based on the titles of the news articles for each season. Common words such as ”US,” ”Stock,” and ”Market” were excluded as they did not effectively represent the event’s topic. The resulting word cloud is shown in Figure 7. It is clear that the economy is mainly influenced by various long-term events. It begins with a gradual decline in the emphasis on COVID. Then, the focus shifted towards managing inflation and the decisions made by the FED. The banking crisis at the start of 2023 soon became the new central point, followed by the acknowledgment of AI as a key driver for the economy, mainly due to the success of LLMs. This indicates that these event trends have the potential to serve as strong predictors of economic indicators and the market. This is also reflected in Appendix D, where we evaluated that news articles have significant predictive power for economic indicators and market trends.
Ticker Frequencies.
{beautifulbox}
Figure 8: The most frequently mentioned stock tickers in the first two years.
Appendix D Additional Results
We test the capacity of our method to explain and predict the excess returns of assets at different levels of performance in the highly complicated market environment in Section D.1. To explore the predictive capability of business news on financial and economic dynamics, we conduct an experiment using refined news features to forecast the economic indicators in Appendix D.2 and market movements in Appendix D.3. We embed the refined news directly and use the daily averaged embeddings of the refined news as predictors in our experiments.
D.1 Decile Portfolios
Figure 9: Cumulative excess return for decile portfolios.
We assess the proposed method in explaining excess returns of assets from different performance levels by applying it to the pricing of decile portfolios. We highlight the first two-year span which shows a highly complicated market environment involving factors like COVID, inflation, the war in Ukraine, and the beginning of the AI explosion as demonstrated in Appendix C. This process includes sorting the assets according to their predicted returns and then forming portfolios for each decile. Figure 9 shows the cumulative excess return over time. The figure clearly demonstrates that each decile creates a distinct ranking of returns in the right position, suggesting that the proposed approach effectively predicts returns at various levels.
D.2 News as Economic Indicators
We assess the capability of news-based features to forecast the daily percentage changes in typical and most popular macroeconomic indicators in different topics sourced from the FRED database [^4] https://fred.stlouisfed.org/. These indicators encompass the stock market (SP500), the market yield on U.S. Treasury Securities at a 10-year constant maturity (DGS10), Moody’s seasoned Baa corporate bond minus the federal funds rate (BAAFF), the 10-year breakeven inflation rate (T10YIE), Brent crude oil prices (DCOILBRENTEU), and the 30-year fixed-rate conforming mortgage index (OBMMIC30YF). The findings are illustrated in Figure 10. The forecasted results exhibit a high degree of accuracy, as evidenced by the high R2 score. This implies that news provides valuable insights for predicting macroeconomic indicators.
D.3 News as Stock Price Predictor
We further investigate the predictive power of news-based features to the price movements of individual stocks. We chose 8 typical stocks that have been frequently mentioned in the news from our analysis in Appendix 8 and used refined news features as predictors to estimate their daily percentage price changes. The results are displayed in Figure 11, with the corresponding R2 scores given in brackets. We note high accuracy and R2 scores for all selected stocks, suggesting that news can significantly help in forecasting stock prices. However, it is important to recognize that due to non-stationarity and the risk of overfitting, stock price prediction cannot be directly applicable as asset pricing Kelly et al. (2023), but it provides insights into the value of the news in the pricing of individual stock.
@@ E.2 Iterative analysis
You are a helpful assistant designed to analyze the business news to assist portfolio management. You will help me analyze this latest news from The Wall Street Journal and provide an analysis report, then I will search the relevant news or articles from the knowledge base based on your analysis report to help you refine iteratively in multiple rounds. Let’s start with this latest news, provide your analysis report, and I will help you refine it with the relevant information later, if you think this news is completely not helpful for investment now or future, call the skip function to skip it, do not skip it if it may contain helpful information to future investment:
@@
Based on your current analysis report, I found potentially relevant news and excerpts from the knowledge base, please refine your analysis report with this information:
@@
Based on your current analysis report, I found potentially relevant news and excerpts from the knowledge base, now finish your analysis report with them:
@@ E.3 Update of note
Now, given the latest news and the analysis report, update the macroeconomic summary with the new insights and impacts from the news. Include any relevant information that could influence the global economic outlook, such as geopolitical events, policy changes, or economic indicators. You should also take note of any important notes about investment trends and chances. Here are the latest news and the analysis report: